### v1

# OmniGuard: An Efficient Approach for AI Safety Moderation Across Modalities

Abstract
The emerging capabilities of large language models (LLMs) have sparked concerns about their immediate potential for harmful misuse. The core approach to mitigate these concerns is the detection of harmful queries to the model. Current detection approaches are fallible, and are particularly susceptible to attacks that exploit mismatched generalization of model capabilities (e.g., prompts in low-resource languages or prompts provided in non-text modalities such as image and audio). To tackle this challenge, we propose OmniGuard, an approach for detecting harmful prompts across languages and modalities. Our approach (i) identifies internal representations of an LLM/MLLM that are aligned across languages or modalities and then (ii) uses them to build a language-agnostic or modality-agnostic classifier for detecting harmful prompts. OmniGuard improves harmful prompt classification accuracy by 11.57% over the strongest baseline in a multilingual setting, by 20.44% for image-based prompts, and sets a new SOTA for audio-based prompts. By repurposing embeddings computed during generation, OmniGuard is also very efficient (120$\approx$ 120$\times$ 120 × faster than the next fastest baseline). Code and data are available at https://github.com/vsahil/OmniGuard.
OmniGuard: An Efficient Approach
for AI Safety Moderation Across Modalities
Sahil Verma1  Keegan Hines2  Jeff Bilmes1  Charlotte Siska2 Luke Zettlemoyer1Hila Gonen1Chandan Singh2 1University of Washington  2Microsoft
1 Introduction
Figure 1: OmniGuard builds a harmfulness classifier that operates on internal representations of an LLM (or MLLM). OmniGuard uses a custom metric (U-Score) to identify representations that generalize across languages and modalities. At inference time, OmniGuard re-uses the embeddings from the LLM/MLLM being used for generation, and thereby completely avoids the overhead of passing the inputs through a separate guard model for safety moderation.
The rapid rise of capabilities in large language models (LLMs) has created an urgent need for safeguards to prevent their immediate harmful misuse as they are deployed to human users en masse Bommasani et al. (2022). Moreover, these safeguards are critical for defending against future potential harms from LLMs Bengio et al. (2024). Standard safeguard approaches broadly include approaches such as safety training using reinforcement learning from human feedback (Ouyang et al., 2022a; Leike et al., 2018) or using pre-trained guard models that classify the safety of an input prompt  (OpenAI, 2025; Inan et al., 2023; Han et al., 2024).
With these safeguards in place, harmful prompts in high-resource languages, e.g., English, are successfully detected. However, harmful prompts in low-resource languages can often bypass these safeguards (Deng et al., 2024; Yong et al., 2024; Yang et al., 2024), i.e., jailbreaking the LLM. Modern LLMs are vulnerable to attacks not only from low-resource natural languages, but also from artificial cipher languages, e.g., base64 or caesar encoding of English prompts (Wei et al., 2023; Yuan et al., 2024a). This phenomenon also extends beyond text to jailbreaking multimodal LLMs (MLLMs) using modalities such as images (Gong et al., 2025; Liu et al., 2024b) or audio (Yang et al., 2025).
Wei et al. (2023) argue that these attacks are successful due to mismatched generalization, a scenario in which the model’s safety training does not generalize to other settings, but general performance does. This might happen since pretraining data often includes more diverse data than that available for safety finetuning (Ghosh et al., 2024b). In this work, we defend against attacks that exploit the mismatched generalization of the safety training of LLMs and MLLMs. Specifically, we defend against attacks that utilize low-resource languages, both natural and cipher languages, as well as attacks employing other modalities, such as images and audio.
We introduce OmniGuard, an approach that builds a classifier using the internal representations of a model. These representations are extracted from specific layers that produce representations that are universally similar across multiple languages and across multiple modalities. OmniGuard’s classifier trained on such representations, is able to accurately detect harmful inputs across 73 languages, with an average of 86.22% accuracy across 53 natural languages and an average of 73.06% accuracy across 20 cipher languages. OmniGuard can also detect harmful inputs provided as images with 88.31% and as audio with 93.09% accuracy respectively.
Compared to popular guard models such as LlamaGuard (Inan et al., 2023), AegisGuard (Ghosh et al., 2024a), or WildGuard (Han et al., 2024) among others, OmniGuard does not require training a separate LLM specifically to detect harmfulness. By building a classifier that uses the internal representations of the main LLM or MLLM, OmniGuard completely avoids the overhead of passing the prompt through a separate guard model, making it very efficient.
In summary, our contributions are the following: (1) We propose OmniGuard, an approach for detecting harmful prompts, (2) we show that OmniGuard accurately detects harmfulness across multiple languages and multiple modalities, (3) we show that OmniGuard is very sample-efficient during training, and (4) we show that OmniGuard is highly efficient at inference time.
2 Methodology
OmniGuard seeks to robustly detect harmful prompts, regardless of their language or modality. We first leverage the tendency of LLMs and MLLMs to create universal representations that are similar across languages (Wendler et al., 2024; Zhao et al., 2024) and across modalities (Wu et al., 2024; Zhuang et al., 2024) in Section 2.1, and then use them to train harmfulness classifiers that robustly detect harmful inputs in Section 2.2.
2.1 Finding language-agnostic representations in an LLM
The first step of OmniGuard searches for internal representations of an LLM that are universally shared across languages. We prompt an LLM with English sentences and their translations to other languages, and extract their representations at different layers. [^1] The representation of a prompt is computed by averaging the representation over each token in the prompt. For language-agnostic representations, we expect the similarity between the representations of English sentences and the representations of their translations to be similar, and we expect this similarity to be higher than the similarity between representations of two sentences that are not translations of each other (a random pair of sentences). We concretize this notion by defining the Universality Score (U-Score, Eq. 1), which is the difference between the average cosine similarities of pairs of sentences that are translations of each other and pairs of sentences that are not.
U-Score ::=:= (1)
$\frac{1}{N}\sum_{\begin{subarray}{c}i\in[N]\end{subarray}}\text{CosSim}\left(\text{Emb}(e_{i}),\text{Emb}(l_{i})\right)$
$-\frac{1}{N(N-1)}$ $\sum_{\begin{subarray}{c}i,j\in[N]\\ i\neq j\end{subarray}}\text{CosSim}\left(\text{Emb}(e_{i}),\text{Emb}(l_{j})\right)$
where $e_{i}$ and $l_{i}$ are sentences in English and their translations to another language.
Figure 2: The U-Score across different layers for different modalities. (A) Different layers of the model Llama3.3-70B-Instruct for different languages. (B) The Cross-Model Alignment Score at different layers of the model (Molmo-7B) for similarity between images and captions. The highest values are obtained with at layers 21-25, indicating better alignment between images and their text captions at these layers. (C) The Cross-Model Alignment Score at different layers of the model (Llama-Omni 8B) for similarity between audios and transcriptions. The highest values are obtained with at layers 20-23, indicating better alignment between audios and their text transcriptions at these layers.
This procedure can be generalized to new different modalities rather than different languages by changing which embeddings are being used. For example, to determine if internal representations of an MLLM are aligned across modalities, we replace embeddings for a translated piece of text with embeddings from a different modality (e.g. a text caption and its corresponding image, or a text transcription and its corresponding audio clip). See experimental details in Section 3.
2.2 Fitting a harmfulness classifier
After selecting the layer that maximizes the U-Score, we extract embeddings from that layer and use them as inputs to fit a lightweight, supervised classifier that predicts harmfulness. In our experiments, the classifier is a multilayer perceptron with 2 hidden layers (with hidden sizes 512 and 256). At inference time, when a prompt is passed to a model for generation, OmniGuard applies this classifier to the embeddings generated by the model, incurring minimal overhead at inference time for safety classification. Note, however, that this approach only applies to open-source models, for which OmniGuard can build a classifier by obtaining embeddings. During training, only the lightweight classifier’s parameters are learned (the original model is never modified), making the training process data-efficient and inexpensive.
3 Experimental Setup
Table 1 and Table 2 give details on all the models and datasets for this section.
Dataset name Citation HuggingFace ID Number of examples
General Flores200 Team et al. (2022) Muennighoff/flores200 997
MM-Vet v2 Yu et al. (2024b) whyu/mm-vet-v2 517
SST-2 (Socher et al., 2013) stanfordnlp/sst2 1000
Text Aegis AI Content Safety Dataset Ghosh et al. (2024b) nvidia/Aegis-AI-Content-Safety-Dataset-1.0 10,800
MultiJail Deng et al. (2024) DAMO-NLP-SG/MultiJail 315
Xsafety Wang et al. (2024a) ToxicityPrompts/XSafety 28,000
RTP-LX de Wynter et al. (2025) ToxicityPrompts/RTP-LX 30,300
AyaRedTeaming Aakanksha et al. (2024) CohereLabs/aya_redteaming 2662
Thai Toxicity tweets Sirihattasak et al. (2018) tmu-nlp/thai_toxicity_tweet 3,300
Ukr Toxicity Dementieva et al. (2024) ukr-detect/ukr-toxicity-dataset 5,000
HarmBench (HB) (Mazeika et al., 2024) walledai/HarmBench 400
Forbidden Questions (FQ) (Shen et al., 2024a) TrustAIRLab/forbidden_question_set 390
Simple Safety Tests (Vidgen et al., 2024) walledai/SimpleSafetyTests 100
SaladBench (SaladB) (Li et al., 2024a) walledai/SaladBench 26,500
Toxicity Jigsaw (TJS) cjadams et al. (2017) Arsive/toxicity_classification_jigsaw 26,000
Toxic Text (Corrêa, 2023) nicholasKluge/toxic-text 41,800
AdvBench (Zou et al., 2023a) walledai/AdvBench 520
CodeAttack (Ren et al., 2024) https://github.com/AI45Lab/CodeAttack 3120
Vision JailBreakV-28K (Luo et al., 2024) JailbreakV-28K/JailBreakV-28k 8,000
VLSafe (Chen et al., 2024c) YangyiYY/LVLM_NLF 1,110
FigStep (Gong et al., 2025) https://github.com/wangyu-ovo/MML 500
MML SafeBench (Wang et al., 2024b) https://github.com/wangyu-ovo/MML 2,510
Hades (Li et al., 2024e) Monosail/HADES 750
SafeBench (Ying et al., 2024) Zonghao2025/safebench 2,300
MM SafetyBench (Liu et al., 2024b) PKU-Alignment/MM-SafetyBench 1680
RedTeamVLM (Li et al., 2024b) MMInstruction/RedTeamingVLM 200
VLSBench (Hu et al., 2025) Foreshhh/vlsbench 2,240
Audio VoiceBench (Alpacaeval) Chen et al. (2024d) hlt-lab/voicebench 636
AIAH (Yang et al., 2025) https://github.com/YangHao97/RedteamAudioLMMs 350
Table 1: Details of datasets used for training and evaluation. Some of the text datasets are inherently multilingual : MultiJail (10 languages), XSafety (10 languages), RTP-LX (28 languages), Aya RedTeaming (8 languages), Thai Toxicity tweets (prompts in Thai), and Ukr Toxicity (prompts in Ukrainian). The remaining text datasets are English-only, and were translated to 72 other languages (52 natural and 20 cipher): HarmBench (HB), Forbidden Questions (FQ), Simple Safety Tests, SaladBench (SaladB), Toxicity Jigsaw (TJS), Toxic Text, and AdvBench.
Model name Citation HuggingFace ID Rough Parameter Count
General Llama3.3-70B-Instruct (Grattafiori et al., 2024) meta-llama/Llama-3.3-70B-Instruct 70B
Molmo-7B (Deitke et al., 2024) allenai/Molmo-7B-D-0924 7B
LLaMA-Omni 8B (Fang et al., 2025) ICTNLP/Llama-3.1-8B-Omni 8B
Kokoro (Hexgrad, 2025) hexgrad/Kokoro-82M 82M
Text LlamaGuard 1 Inan et al. (2023) meta-llama/LlamaGuard-7b 7B
LlamaGuard 2 Inan et al. (2023) meta-llama/Meta-Llama-Guard-2-8B 8B
LlamaGuard 3 Inan et al. (2023) meta-llama/Llama-Guard-3-8B 8B
AegisGuard Permissive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Permissive-1.0 7B
AegisGuard Defensive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Defensive-1.0 7B
WildGuard Han et al. (2024) allenai/wildguard 7B
HarmBench (mistral) Mazeika et al. (2024) cais/HarmBench-Mistral-7b-val-cls 7B
HarmBench (llama) Mazeika et al. (2024) cais/HarmBench-Llama-2-13b-cls 13B
DuoGuard Deng et al. (2025) DuoGuard/DuoGuard-1B-Llama-3.2-transfer 1B
PolyGuard Kumar et al. (2025) ToxicityPrompts/PolyGuard-Qwen 7B
Vision Llama Guard 3 Vision (Chi et al., 2024) meta-llama/Llama-Guard-3-11B-Vision 11B
VLMGuard (Du et al., 2024) --- 2.2M
LLavaGuard (Helff et al., 2025) AIML-TUDA/LlavaGuard-7B-hf 7B
Table 2: Model and baseline details.
3.1 Selecting universal layers via the U-Score
Selecting language-agnostic layers
To select language-agnostic layers, we use a dataset of translated sentence pairs spanning various languages. Specifically, we use sentences in 53 natural languages from the Flores200 dataset and additionally translate the sentences into 20 cipher languages (using encodings such as Caesar shifts, base64, hexadecimal); see a full list in Appendix A. We extract embeddings from each layer of Llama3.3-70B-Instruct for the sentences in all 73 languages and use them to compute the U-Score (averaged over languages). Fig. 2 shows the U-Score as a function of layer depth. For natural languages (blue curve), the U-Score peaks in the middle layers of the model, with the highest values in layer 57 (out of 81 layers). For cipher languages (red curve), the U-Score is much lower than for natural languages, suggesting the model fails to represent semantic similarity in these languages (see analysis in Section 5).
Selecting modality-agnostic layers
To select layers aligned between images and captions, we use the MM-Vet v2 dataset, a popular dataset for MLLM evaluation containing 517 examples, each consisting of a text question paired with one or more images. We generate captions for each image using a captioning model (Molmo-7B) and then extract embeddings for each image and its corresponding caption using an MLLM (also Molmo-7B) and use them to compute the U-Score, which peaks in layer 22 (out of 28 layers; see Fig. 2 green curve).
To select layers aligned between text and audio, we use the audio version of the Alpacaeval dataset from VoiceBench, a dataset of 636 audio-transcript pairs. We extract embeddings from each layer of an MLLM (LLaMA-Omni 8B) and use them to compute the U-Score, which peaks at layer 21 (out of 32 layers; see Fig. 2 purple curve).
Overall, we see that LLMs and MLLMs generate representations that are shared across languages and modalities.
3.2 Training and evaluating the harmfulness classifier
3.2.1 Setup for multilingual text attacks
OmniGuard classifier.
Following Section 3.1, we build a classifier that takes as input embeddings from layer 57 of Llama3.3-70B-Instruct. As training data, we randomly select 2,800 examples from the Aegis AI Content Safety dataset, balancing the benign and harmful classes. Notably, this dataset is about 18$\times$ smaller than the training data used by our baseline methods. We translate these English examples to 52 other natural languages (via the Google Translate API) and 20 cipher languages (using fixed rules), totaling 73 languages. We train OmniGuard using only half the languages (see list in Appendix A).
Baselines.
We compare to many popular guard models (see Table 2) middle row. Notably, DuoGuard and PolyGuard were trained to detect harmful prompts across multiple languages. For a more direct comparison, we also compare to finetuned versions of DuoGuard and PolyGuard using the same 37 languages we use to train OmniGuard; Following the original PolyGuard paper, we finetuned these models using LoRA (Hu et al., 2021) for all linear layers with rank 8 and alpha 16 for one epoch with a learning rate of $42e-$ 4.
Datasets.
We evaluate on several common text attack benchmarks (see Table 1). We additionally evaluate on three benchmarks from CodeAttacks that transform a harmful query as a list, a stack, or as a string in a Python program, obfuscating the harmfulness. For evaluation in this setup, we transform the harmful prompts from AdvBench and benign prompts from Toxicity Jigsaw datasets in the three code formats and subsampl the Toxicity Jigsaw dataset to be of the same size as Advbench. Note that, for this experiment, we only trained OmniGuard on the English subset of the training dataset.
3.2.2 Setup for vision attacks
OmniGuard classifier.
Following Section 3.1, we build a classifier that takes as input embeddings from layer 22 of Molmo-7B. As training data, we use 2000 image-query pairs randomly sampled from the JailBreakV-28K dataset and 1024 image-query pairs sampled from the VLSafe dataset as the harmful datapoints and 517 image-query from the MM-Vet v2 dataset as the benign datapoints.
Baselines.
We compare to guard models that take an image or image-text pair and output a binary harmfulness classification (see Table 2 bottom row). We train VLMGuard on the same training data as OmniGuard.
Datasets.
We evaluate detecting image/text attacks using several datasets (see Table 1). FigStep and MML Safebench are typographic attacks that embed a harmful prompt in an image. MML Safebench further encrypts a harmful prompt in several variants, such as rotation, mirror images, word replacement, and with base64 encoding. Hades and Safebench consist of images and text queries where the text itself is harmful. MM-safetybench, RTVLM, and VLSBench consist of an image and a query where the text query is seemingly benign, but when combined with the respective image, it is harmful (e.g. see Figure 1).
3.2.3 Setup for audio attacks
OmniGuard classifier.
Following Section 3.1, we build a classifier that takes as input embeddings from layer 21 of Llama-Omni-8B. We train the classifier on the English portion of the training data we use for the text setting, by using a text-to-speech model to convert the text into audio. We use the open-source Kokoro model as the text-to-speech model.
Baselines.
We are unaware of any existing models for detecting harmful audio input. The most relevant approach, SpeechGuard (Peri et al., 2024) adds noise as a defense against potentially harmful audio inputs but does not directly classify harmfulness. To contextualize our results for audio benchmarks, we compare performance to guard models that directly classify the raw text present in the audio (OmniGuard and LlamaGuard3).
Datasets.
We use the two audio benchmarks (see Table 1 bottom row). We also evaluate on several text jailbreak benchmarks using Kokoro to convert them from text to speech: HB, FQ, Simple Safety Tests, SaladB, and TJS. We use Kokoro for generating text-to-speech versions.
4 Results
MultiJail Xsafety RTP-LX Aya RedTeaming Thai Tox Ukr Tox Avg. (A) Multilingual text benchmarks LlamaGuard 1 39.27 57.01 48.66 54.49 41.31 53.99 49.12 LlamaGuard 2 48.69 52.66 34.69 58.58 42.86 51.79 48.21 LlamaGuard 3 66.87 64.34 45.57 63.83 46.73 51.79 56.52 AegisGuard (P) 61.49 79.78 75.07 78.88 56.09 65.75 69.51 AegisGuard (D) 79.71 90.77 92.17 89.78 63.34 67.95 80.62 WildGuard 42.55 71.23 71.94 61.45 40.42 55.03 57.10 HarmBench (llama) 0.22 0.14 0.0 0.03 39.04 50.1 14.92 HarmBench (mistral) 2.4 5.65 5.14 7.39 40.42 50.55 18.59 MD-Judge 25.78 53.58 66.46 46.20 39.48 53.89 47.56 DuoGuard 39.20 63.42 66.57 61.80 45.63 50.75 54.56 PolyGuard 82.00 96.41 83.86 90.34 70.43 76.07 83.19 OmniGuard 93.83 93.64 94.55 94.31 68.7 73.1 86.36 HarmBench FQ SimpleST SaladB TJS ToxText AdvBench Avg. (B) Translated text benchmarks LlamaGuard 1 32.47 23.75 34.32 23.27 62.49 65.55 34.39 39.46 LlamaGuard 2 57.19 43.72 50.71 34.54 58.21 62.17 56.95 51.93 LlamaGuard 3 70.02 53.25 67.81 46.30 62.33 70.87 70.26 62.98 AegisGuard (P) 62.16 43.01 56.55 44.92 73.69 72.80 62.12 59.32 AegisGuard (D) 88.53 76.67 87.64 78.27 71.38 68.72 90.77 80.28 WildGuard 33.64 31.20 33.90 27.37 66.61 67.27 39.98 42.85 HarmBench (llama) 0.03 0.11 0.01 0.07 48.62 49.97 0.01 14.12 HarmBench (mistral) 2.32 1.75 2.04 1.66 50.53 50.69 1.7 15.81 MD-Judge 16.19 12.11 22.29 13.81 65.34 64.26 25.67 31.38 DuoGuard 20.44 44.36 28.79 36.88 68.57 69.07 28.58 42.38 PolyGuard 66.22 56.05 62.53 54.88 78.34 76.52 67.96 66.07 OmniGuard 89.13 89.57 89.62 87.30 76.68 75.07 86.59 84.85 HarmBench FQ SimpleST SaladB TJS ToxText AdvBench Avg. (C) Unseen langs. FT DuoGuard 23.59 39.08 28.14 33.29 54.1 53.23 28.29 37.1 FT PolyGuard 72.45 79.84 76.81 76.85 74.07 72.33 73.55 75.13 OmniGuard 86.51 86.65 86.42 85.01 72.82 71.44 84.29 81.88
Table 3: Accuracy of detecting harmful prompts for text attack benchmarks that are (A) multilingual benchmarks, (B) English translated to 73 languages, and (C) English translated to languages not seen at training time. In all settings, OmniGuard achieves the highest performance. Table B1 further stratifies these results by high-resource, low-resource, and cipher languages.
Hades VLSBench MM-SafetyBench SafeBench RTVLM FigStep Avg. (A) Image +Query Llama3 Vision GRD 76.00 3.97 31.90 68.40 56.50 47.40 47.36 VLMGuard 98.00 74.56 92.20 73.90 94.00 99.80 88.74 LLavaGuard 23.73 42.08 10.95 12.10 18.50 3.40 18.46 OmniGuard 100.00 92.24 99.82 91.60 89.00 100.00 95.44 MML Rotate MML Mirror MML W.R. MML Q.R. MML Base64 Avg. (B) Typographed image Llama3 Vision GRD 83.20 68.00 96.40 25.40 98.80 74.36 VLMGuard 6.80 21.00 100.0 86.20 0.20 42.84 LLavaGuard 0.00 0.00 0.00 11.40 0.00 2.28 OmniGuard 100.0 100.0 99.60 98.80 0.40 79.76
Table 4: Accuracy of detecting harmful queries in multimodal benchmarks for (A) image-query pairs and (B) typographed images with encrypted text. OmniGuard achieves the highest performance for both kinds of benchmarks.
AIAH SafeBench (M) SafeBench (F) HB FQ SimpleST SaladB TJS AdvBench OmniGuard (Audio) 91.14 94.4 93.8 95.98 90.42 97.0 94.21 82.03 98.85 OmniGuard (text-en) - - - 92.0 93.3 93.0 90.2 93.2 90.0 LlamaGuard3 (text-en) - - - 97.32 78.75 99.0 67.03 72.16 98.07
Table 5: Accuracy of detecting harmful queries in audio. OmniGuard is able to detect harmful audio inputs with high accuracy across all benchmarks. Since there are no baselines for detecting harmful prompts in audio, we compare the performance against OmniGuard’s and LlamaGuard3 when the same benchmarks are provided as text in English.
Defending against multilingual text attacks
Table 3 compares the accuracy of detecting harmful prompts for text benchmarks. Table 3(A) shows results for multilingual benchmarks, where OmniGuard achieves the highest accuracy (86.36%) compared to the baselines, and achieves new state-of-the-art performance for 3 benchmarks: MultiJail, RTP-LX, and AyaRedTeaming. The strongest baseline is Polyguard, which yields an average accuracy of 83.19%, despite being trained on a much larger dataset (1.91M examples for Polyguard versus 103K examples for OmniGuard). In benchmarks that were translated from English to various other languages, including cipher languages, we again see that OmniGuard achieves the highest accuracy (Table 3(B)). Finally, Table 3(C) shows that OmniGuard outperforms finetuned versions of DuoGuard and Polyguard on unseen languages, demonstrating that OmniGuard can outperform methods that were trained specifically for multilingual harmfulness classification.
Defending against image-based attacks
Table 4 shows the accuracy of detecting harmful image and text prompts for (A) pairs consisting of images and text queries, where either the image or both the image and query can be harmful and (B) typographic images with various encryptions. OmniGuard achieves the highest performance for both sets of benchmarks (95.44% and 79.76%) while being trained using only about 3500 image-query pairs (compared to about 5500 datapoints used by LlavaGuard). The only benchmark where OmniGuard fails to detect harmful prompts is MML Base64, which consists of typographed images of prompts encrypted using base64 encoding.
Defending against audio-based attacks
Table 5 shows the accuracy of detecting harmful audio prompts. OmniGuard detects harmful audio input with high accuracy across all benchmarks. As we are not aware of any existing defenses for audio jailbreaks, we compare against OmniGuard and LlamaGuard3’s accuracy in detecting harmful prompts when the same inputs are provided in English text. The accuracy OmniGuard achieves in detecting harmful audio inputs is similar to or higher than its performance for detecting harmful text inputs.
Figure 3: Accuracy of detecting harmful prompts in a few-shot setting. As few-shot examples are provided, OmniGuard quickly achieves near-perfect accuracy, despite the attacks being quite different from its training data (e.g. without any few-shot examples, OmniGuard’s accuracy is close to 50% ). In contrast, the guard model baselines improve their accuracy slowly in a few-shot setting, despite sometimes having seen similar code attacks in their training data. Accuracies are averaged over 50 random sets of few-shot examples; error bars show the standard error of the mean.
Data-efficient adaptation
We also evaluate the accuracy of OmniGuard and baselines in adapting to out-of-distribution code attacks given very few samples. In this setting, some prior work has speculated that guard models may be very data efficient, as they can make use of few-shot examples in-context (Inan et al., 2023). However, we find that baseline guard models generally struggle to rapidly adapt to this setting given few-shot examples (Figure 3). [^2] Note, we omit baseline guard models that achieve 90% accuracy or greater without any few-shot examples, as their training data likely explicitly includes code attacks. In contrast, OmniGuard is able to rapidly achieve close to 100% accuracy for all three benchmarks by updating its lightweight parameters using less than five examples.
5 Analysis
Effect of U-Score-based layer selection.
We perform ablation experiments to determine the effect of selecting the appropriate layer for training the OmniGuard classifier. For the text-only model, we compare the U-Score-selected layer (57) to 3 other layers (layer 10, layer 75, and the last layer) when used for a set of toxicity prediction tasks. Table 6 shows that the representations from the layer with the highest U-Score result in significantly better harmfulness classification accuracy, improving between 5% and 14% compared to the other layers.
Thai Tox Ukr Tox TJS ToxText Avg.
Layer 10 62.1 65.5 66.95 61.89 64.42
Layer 75 65.2 66.4 70.72 65.79 68.26
Last Layer 63.1 51.2 61.33 56.76 59.05
U-Score selected layer (57) 68.7 73.1 76.8 75.07 73.4
Table 6: OmniGuard’s accuracy of detecting harmful prompt when trained using representation from different model layers.
OmniGuard’s efficiency
OmniGuard is highly efficient at inference time because it re-uses the internal representations of the main LLM that is already processing the user query for generation. Therefore, its compute time is only that of a lightweight multilayer perceptron, making it much faster than baseline guard models (note that this does limit OmniGuard to only work when the generation model is open-source, so embeddings can be extracted). Table 7 shows the inference time required by various guard models to predict the harmfulness of prompts in the AdvBench dataset in English, translated to Spanish, French, Telugu, and base64 encoding. OmniGuard is the fastest and is about 120$\times$ faster than the fastest baseline (DuoGuard). Inference time as measured on a machine with 1 L40 GPU, 4 CPUs, and 50 GB RAM.
Guard Method Inference Time (s) $\downarrow$
LlamaGuard 3 87.25
AegisGuard (D) 152.26
WildGuard 306.14
MD-Judge 128.26
DuoGuard 4.85
PolyGuard 409.90
OmniGuard 0.04
Table 7: Average inference time required for harmfulness prediction on the AdvBench dataset (averaged over 5 languages). OmniGuard is about 120$\times$ faster than the fastest baseline (DuoGuard).
Performance comparison across languages.
We now analyze the harmfulness classification accuracy of OmniGuard by language, and compare it to the underlying LLM’s sentiment classification accuracy for the same language (Fig. 4). We measure harmfulness classification accuracy using OmniGuard on all the datasets in Table 3 and sentiment classification accuracy using Llama3.3-70B-Instruct with zero-shot prompting on 72 translated versions of the SST-2 dataset (translated to all the languages we consider).
We observe that the accuracies are generally correlated indicating that OmniGuard is able to defend well for languages for which the LLM is more coherent / susceptible to attack. Unsurprisingly, the accuracies for natural languages are higher than the accuracies for cipher languages. Nevertheless, harmfulness classification accuracy can be fairly high, even when sentiment classification accuracy is near chance (50%).
Figure 4: Comparison of accuracy of classifying sentiments in various languages compared to detecting harmful prompts in those languages using OmniGuard. In both cases the LLM is Llama3.3-70B-Instruct.
6 Related Work
Jailbreak Attacks in LLMs
Several techniques have recently emerged to attack or jailbreak LLMs. Early techniques relied on manual effort and were very time-intensive Shen et al. (2024b); Andriushchenko et al. (2024). Later techniques automated this process, e.g., Zou et al. (2023b); Jones et al. (2023); Zhu et al. (2023) proposed gradient-based approaches to identify inputs to jailbreak LLMs with white-box access. Another set of techniques start from a set of human written prompts and modify them using approaches like genetic algorithms (Liu et al., 2024a; Lapid et al., 2024; Li et al., 2024d), fuzzing (Yu et al., 2024a), or reinforcement learning (Chen et al., 2024b) to automatically produce prompts for jailbreaking. Another set of techniques, use a helper LLM to generate prompts that attack a target LLM (Chao et al., 2024; Ding et al., 2024; Mehrotra et al., 2024). Finally, Wei et al. (2024); Wang et al. (2023); Anil et al. (2024); Pernisi et al. (2024) use simple in-context demonstrations to jailbreak the models by overcoming its safety training and Russinovich et al. (2024); Li et al. (2024c) propose using multi-turn dialogues to jailbreak models.
Multilingual Jailbreak Attacks
Most of the aforementioned jailbreak techniques focus on attacks in English, against which significant defense exists both at the model and system level. To tackle this, a novel set of techniques have emerged that attack models using inputs in various languages or obfuscations that are able to bypass the safety guardrails. Deng et al. (2024); Yong et al. (2024); Wang et al. (2024a); Yang et al. (2024); Yoo et al. (2024); Upadhayay and Behzadan (2024); Song et al. (2024) demonstrated that attacking models using mid and low resources languages led to higher attack success rates, compared to the case of attacking the model in high-resource languages like English.
Going beyond natural languages, a newer set of works propose using cipher characters or languages to evade the safety filters, e.g., Jin et al. (2024) propose interspersing cipher characters in between text, Jiang et al. (2024) propose replacing the unsafe words with their ASCII art versions, and Yuan et al. (2024a) propose prompting models in cipher languages like Morse, Atbash, Caesar.
Multimodal Jailbreak Attacks
Using modalities apart from text aims to explore a completely new attack surface, like images or audios. Several recent works have shown that MLLMs remain vulnerable to being jailbroken when prompted with images or audios that have a harmful query (the same harmful query in text would be easily detected as harmful). Liu et al. (2024b) show that using a prompt with a correlated image, e.g., using an image of a bomb when asking the model to answer the question: How to make a bomb? is more likely to jailbreak a model than when using an uncorrelated image. Hu et al. (2025) argue that providing a harmful image with a benign query (see Figure 1) further increases the potential of jailbreaking the model. Gong et al. (2025) and Wang et al. (2024b) demonstrate jailbreaking models by simply typographically embedding harmful queries in an image.
Safety moderation in LLMs
Safety moderation in LLMs broadly fits into two categories: intrinsic and extrinsic. Intrinsic mechanisms include finetuning or RLHF training on an LLM (Bianchi et al., 2024; Chen et al., 2024a; Yuan et al., 2024b; Ouyang et al., 2022b; Bai et al., 2022; Dai et al., 2023). Extrinsic safety mechanisms utilize external models to detect harmful inputs and responses; these models can either be simple filters or use guard models. Jain et al. (2023); Alon and Kamfonas (2023); Hu et al. (2024) propose using perplexity filtering for detecting harmful prompts. Guard models defend LLMs by training separate LLMs to detect harmful text (see Table 2).
Safety against multilingual attacks
DuoGuard (Deng et al., 2025) and PolyGuard (Kumar et al., 2025) are the two previous guard models that were specifically trained to defend against multilingual attacks. DuoGuard uses a two-player RL-driven mechanism to generate harmful data in multiple languages and uses that to finetune a Llama3.2-1B model. PolyGuard collects an extensive dataset of 1.91M samples of harmful and benign datapoints in 17 languages and uses that to finetune a Qwen-2.5-7B-Instruct model.
Safety moderation in MLLMs
Relatively few works have tackled detecting harmful prompts in multimodal settings (see Table 1 and Table 2). Chi et al. (2024) propose LlamaGuard3-11B-Vision (a finetuned version of Llama-3-11B-Vision) for detecting unsafe inputs in images and texts. Du et al. (2024) and Helff et al. (2025) propose other approaches for the same task. OmniGuard achieves higher accuracy in detecting harmful images and prompts compared to these approaches, and to the best of our knowledge is the first guard model for harmful audio inputs.
7 Conclusions
We propose OmniGuard, an approach for training a safety moderation classifier using the internal representations of an LLM or MLLM that are universally similar across languages and modalities. Our approach consists of two steps: first, we identify these universally similar representations and then we use them to train a harmfulness classifier. We find that OmniGuard accurately detects harmful prompts across languages, including low-resource languages as well as cipher languages, and also across modalities – images and audios. We show that OmniGuard allows to train more efficient safety moderation classifiers (both in training time and in inference time) compared to standard guard models, and conclude that our approach is superior in both accuracy and efficiency across languages and modalities.
Limitations
While OmniGuard achieves state-of-the-art performance for detecting harmful prompts across languages and modalities, its performance depends on the underlying model. If the underlying model does not understand the language or an image or audio input, OmniGuard might not be able to detect if the input is harmful. However, this limitation is not unique to OmniGuard, and existing approaches suffer from the same limitation.
Our approach also relies on the existence of universally similar representations, which we empirically found to exist across models and modalities. However, we did not exhaustively check all models and this assumption might not hold for models that we have not used in this work. Moreover, OmniGuard requires access to internal representations of a model, making it inapplicable to closed-source models.
Lastly, the results we report are based on a fixed set of evaluation datasets that are standard benchmarks used in the research area of AI safety moderation. While OmniGuard performs well across the datasets we experiment with, its performance in real-world settings might differ.
Ethics. While this work seeks to mitigate the risks of LLM deployment in high-risk scenarios, OmniGuard is not a perfect classifier and unexpected failures may allow for the harmful misuse of LLMs.
Appendix A Languages Used in Our Approach
We use the following languages in our experiments:
  1. 1.
    Natural Languages: English, French, German, Spanish, Persian, Arabic, Croatian, Japanese, Polish, Russian, Swedish, Thai, Hindi, Italian, Korean, Bengali, Portuguese, Chinese, Hebrew, Serbian, Danish, Turkish, Greek, Indonesian, Zulu, Hungarian, Basque, Swahili, Afrikaans, Bosnian, Lao, Romanian, Slovenian, Ukrainian, Finnish, Malay, Javanese, Welsh, Bulgarian, Armenian, Icelandic, Vietnamese, Sinhalese, Maori, Gujarati, Kannada, Marathi, Tamil, Telugu, Amharic, Norwegian, Czech, Dutch.
  2. 2.
    Cipher Languages: Caesar1, Caesar2, Caesar3, Caesar4, Caesar5, Caesar6, Caesar7, Caesarneg1, Caesarneg2, Caesarneg3, Caesarneg4, Caesarneg5, Caesarneg6, Caesarneg7, Ascii, Hexadecimal, Base64, Leet, Vowel, Alphanumeric. A number in front of Caesar cipher means that the English alphabets were shifted by that much forward and a number in front of Caesarneg cipher means that the English alphabets were shifted by that much backward.
Out of these languages, we use the following for training our classifier: Arabic, Chinese, Czech, Dutch, English, French, German, Hindi, Italian, Japanese, Korean, Polish, Portuguese, Russian, Spanish, Swedish, Thai, Bosnian, Turkish, Finnish, Indonesian, Bengali, Swahili, Vietnamese, Tamil, Telugu, Greek, Maori, Javanese, Caesar1, Caesar2, Caesar4, Caesarneg2, Caesarneg4, Caesarneg6, Ascii, Hexadecimal
And these for testing: Persian, Croatian, Hebrew, Serbian, Danish, Zulu, Hungarian, Basque, Afrikaans, Lao, Romanian, Slovenian, Ukrainian, Malay, Welsh, Bulgarian, Armenian, Icelandic, Sinhalese, Gujarati, Kannada, Marathi, Amharic, Norwegian, Caesar, Caesar5, Caesar7, Caesarneg3, Caesarneg1, Caesar6, Caesarneg7, Caesarneg5, Base64, Alphanumeric, Vowel, LeetSpeak.
Appendix B Datasets and models
High-Res Low-Res Cipher LlamaGuard 1 69.92 41.25 16.07 LlamaGuard 2 75.28 62.2 16.2 LlamaGuard 3 82.23 75.84 24.74 AegisGuard (P) 83.36 59.06 44.22 AegisGuard (D) 88.14 76.26 83.21 WildGuard 81.35 43.51 16.53 HarmBench (llama) 14.25 14.08 14.11 HarmBench (mistral) 17.6 15.9 14.46 MD-Judge 59.51 30.21 15.44 DuoGuard 71.4 46.26 15.77 PolyGuard 94.47 79.22 21.28 OmniGuard 88.25 85.56 73.06
Table B1: Accuracy of detecting harmful prompts stratified by high-resource natural, low-resource natural, and cipher languages.

### v2

@@
-OmniGuard: An Efficient Approach for AI Safety Moderation Across Modalities
+OmniGuard: An Efficient Approach for AI Safety Moderation Across Languages and Modalities
 Abstract
@@
 OmniGuard improves harmful prompt classification accuracy by 11.57% over the strongest baseline in a multilingual setting, by 20.44% for image-based prompts, and sets a new SOTA for audio-based prompts.
-By repurposing embeddings computed during generation, OmniGuard is also very efficient (120$\approx$ 120$\times$ 120 × faster than the next fastest baseline).
+By repurposing embeddings computed during generation, OmniGuard is also very efficient ($\approx\!120\times$ faster than the next fastest baseline).
 Code and data are available at https://github.com/vsahil/OmniGuard.
 OmniGuard: An Efficient Approach
-for AI Safety Moderation Across Modalities
+for AI Safety Moderation Across Languages and Modalities
 Sahil Verma1   Keegan Hines2   Jeff Bilmes1   Charlotte Siska2 Luke Zettlemoyer1Hila Gonen1Chandan Singh2 1University of Washington  2Microsoft
@@
 At inference time, OmniGuard re-uses the embeddings from the LLM/MLLM being used for generation, and thereby completely avoids the overhead of passing the inputs through a separate guard model for safety moderation.
-The rapid rise of capabilities in large language models (LLMs) has created an urgent need for safeguards to prevent their immediate harmful misuse as they are deployed to human users en masse Bommasani et al.
-(2022).
-Moreover, these safeguards are critical for defending against future potential harms from LLMs Bengio et al.
-(2024).
-Standard safeguard approaches broadly include approaches such as safety training using reinforcement learning from human feedback (Ouyang et al., 2022a; Leike et al., 2018) or using pre-trained guard models that classify the safety of an input prompt  (OpenAI, 2025; Inan et al., 2023; Han et al., 2024).
+The rapid rise of capabilities in large language models (LLMs) has created an urgent need for safeguards to prevent their immediate harmful misuse as they are deployed to human users en masse bommasani2022-opportunities-risks.
+Moreover, these safeguards are critical for defending against future potential harms from LLMs bengio2024managing.
+Standard safeguard approaches broadly include approaches such as safety training using reinforcement learning from human feedback (instruct-gpt; leike2018-scalable-agent-alignment) or using pre-trained guard models that classify the safety of an input prompt  (openai2025moderation; inan2023-llama-guardrails; Han2024-wildguardmix-dataset).
 With these safeguards in place, harmful prompts in high-resource languages, e.g., English, are successfully detected.
-However, harmful prompts in low-resource languages can often bypass these safeguards (Deng et al., 2024; Yong et al., 2024; Yang et al., 2024), i.e., jailbreaking the LLM.
-Modern LLMs are vulnerable to attacks not only from low-resource natural languages, but also from artificial cipher languages, e.g., base64 or caesar encoding of English prompts (Wei et al., 2023; Yuan et al., 2024a).
-This phenomenon also extends beyond text to jailbreaking multimodal LLMs (MLLMs) using modalities such as images (Gong et al., 2025; Liu et al., 2024b) or audio (Yang et al., 2025).
-Wei et al.
-(2023) argue that these attacks are successful due to mismatched generalization, a scenario in which the model’s safety training does not generalize to other settings, but general performance does.
-This might happen since pretraining data often includes more diverse data than that available for safety finetuning (Ghosh et al., 2024b).
+However, harmful prompts in low-resource languages can often bypass these safeguards (deng2024-multilingual-attack-firstpaper; yong2024-low-resource-languages-jailbreak-gpt4; dan-roth-paper-guard-llms-on-multilingual-data), i.e., jailbreaking the LLM.
+Modern LLMs are vulnerable to attacks not only from low-resource natural languages, but also from artificial cipher languages, e.g., base64 or caesar encoding of English prompts (jacob-steinhardt-why-jailbroken; yuan2024-gpt-cipher-attack).
+This phenomenon also extends beyond text to jailbreaking multimodal LLMs (MLLMs) using modalities such as images (figstep-dataset-image-jailbreak; mm-safetybench-image-jailbreak) or audio (achilles-dataset-audio-jailbreaks).
+jacob-steinhardt-why-jailbroken argue that these attacks are successful due to mismatched generalization, a scenario in which the model’s safety training does not generalize to other settings, but general performance does.
+This may happen because pretraining data often includes more diverse data than that available for safety finetuning (ghosh2024aegis2).
 In this work, we defend against attacks that exploit the mismatched generalization of the safety training of LLMs and MLLMs.
@@
 OmniGuard can also detect harmful inputs provided as images with 88.31% and as audio with 93.09% accuracy respectively.
-Compared to popular guard models such as LlamaGuard (Inan et al., 2023), AegisGuard (Ghosh et al., 2024a), or WildGuard (Han et al., 2024) among others, OmniGuard does not require training a separate LLM specifically to detect harmfulness.
-By building a classifier that uses the internal representations of the main LLM or MLLM, OmniGuard completely avoids the overhead of passing the prompt through a separate guard model, making it very efficient.
+In contrast to popular guard models such as LlamaGuard (inan2023-llama-guardrails), AegisGuard (ghosh2024-aegis-safety-guard), or WildGuard (Han2024-wildguardmix-dataset), OmniGuard does not require training a separate LLM specifically to detect harmfulness.
+By building a classifier that uses the internal representations of the main LLM or MLLM, OmniGuard avoids the overhead of passing the prompt through a separate guard model, making it very efficient.
 In summary, our contributions are the following: (1) We propose OmniGuard, an approach for detecting harmful prompts, (2) we show that OmniGuard accurately detects harmfulness across multiple languages and multiple modalities, (3) we show that OmniGuard is very sample-efficient during training, and (4) we show that OmniGuard is highly efficient at inference time.
@@
 OmniGuard seeks to robustly detect harmful prompts, regardless of their language or modality.
-We first leverage the tendency of LLMs and MLLMs to create universal representations that are similar across languages (Wendler et al., 2024; Zhao et al., 2024) and across modalities (Wu et al., 2024; Zhuang et al., 2024) in Section 2.1, and then use them to train harmfulness classifiers that robustly detect harmful inputs in Section 2.2.
+We first leverage the tendency of LLMs and MLLMs to create universal representations that are similar across languages (wendler2024-convert-things-to-english; zhao2024-convert-things-to-english) and across modalities (wu2024-semantic-hub-hypothesis; zhuang2024vector) in Section˜2.1, and then use them to train harmfulness classifiers that robustly detect harmful inputs in Section˜2.2.
 2.1 Finding language-agnostic representations in an LLM
@@
 1), which is the difference between the average cosine similarities of pairs of sentences that are translations of each other and pairs of sentences that are not.
-U-Score ::=:= (1)
+U-Score = (1)
 $\frac{1}{N}\sum_{\begin{subarray}{c}i\in[N]\end{subarray}}\text{CosSim}\left(\text{Emb}(e_{i}),\text{Emb}(l_{i})\right)$
@@
 a text caption and its corresponding image, or a text transcription and its corresponding audio clip).
-See experimental details in Section 3.
+See experimental details in Section˜3.
 2.2 Fitting a harmfulness classifier
@@
 3 Experimental Setup
-Table 1 and Table 2 give details on all the models and datasets for this section.
+Table˜1 and Table˜2 give details on all the models and datasets for this section.
 Dataset name Citation HuggingFace ID Number of examples
-General Flores200 Team et al. (2022) Muennighoff/flores200 997
-MM-Vet v2 Yu et al. (2024b) whyu/mm-vet-v2 517
-SST-2 (Socher et al., 2013) stanfordnlp/sst2 1000
-Text Aegis AI Content Safety Dataset Ghosh et al. (2024b) nvidia/Aegis-AI-Content-Safety-Dataset-1.0 10,800
-MultiJail Deng et al. (2024) DAMO-NLP-SG/MultiJail 315
-Xsafety Wang et al. (2024a) ToxicityPrompts/XSafety 28,000
-RTP-LX de Wynter et al. (2025) ToxicityPrompts/RTP-LX 30,300
-AyaRedTeaming Aakanksha et al. (2024) CohereLabs/aya_redteaming 2662
-Thai Toxicity tweets Sirihattasak et al. (2018) tmu-nlp/thai_toxicity_tweet 3,300
-Ukr Toxicity Dementieva et al. (2024) ukr-detect/ukr-toxicity-dataset 5,000
-HarmBench (HB) (Mazeika et al., 2024) walledai/HarmBench 400
-Forbidden Questions (FQ) (Shen et al., 2024a) TrustAIRLab/forbidden_question_set 390
-Simple Safety Tests (Vidgen et al., 2024) walledai/SimpleSafetyTests 100
-SaladBench (SaladB) (Li et al., 2024a) walledai/SaladBench 26,500
-Toxicity Jigsaw (TJS) cjadams et al. (2017) Arsive/toxicity_classification_jigsaw 26,000
-Toxic Text (Corrêa, 2023) nicholasKluge/toxic-text 41,800
-AdvBench (Zou et al., 2023a) walledai/AdvBench 520
-CodeAttack (Ren et al., 2024) https://github.com/AI45Lab/CodeAttack 3120
-Vision JailBreakV-28K (Luo et al., 2024) JailbreakV-28K/JailBreakV-28k 8,000
-VLSafe (Chen et al., 2024c) YangyiYY/LVLM_NLF 1,110
-FigStep (Gong et al., 2025) https://github.com/wangyu-ovo/MML 500
-MML SafeBench (Wang et al., 2024b) https://github.com/wangyu-ovo/MML 2,510
-Hades (Li et al., 2024e) Monosail/HADES 750
-SafeBench (Ying et al., 2024) Zonghao2025/safebench 2,300
-MM SafetyBench (Liu et al., 2024b) PKU-Alignment/MM-SafetyBench 1680
-RedTeamVLM (Li et al., 2024b) MMInstruction/RedTeamingVLM 200
-VLSBench (Hu et al., 2025) Foreshhh/vlsbench 2,240
-Audio VoiceBench (Alpacaeval) Chen et al. (2024d) hlt-lab/voicebench 636
-AIAH (Yang et al., 2025) https://github.com/YangHao97/RedteamAudioLMMs 350
+General Flores200 nllbteam2022-languageleftbehindscaling Muennighoff/flores200 997
+MM-Vet v2 yu2024-mmvetv2-dataset whyu/$m$-vet-v2 517
+SST-2 (socher-etal-SSt-5-dataset) stanfordnlp/sst2 1000
+Text Aegis AI Content Safety Dataset ghosh2024aegis2 nvidia/Aegis-AI-Content-Safety-Dataset-1.0 10,800
+MultiJail deng2024-multilingual-attack-firstpaper DAMO-NLP-SG/MultiJail 315
+Xsafety wang2024-multilingualsafety ToxicityPrompts/XSafety 28,000
+RTP-LX rtplx-dataset ToxicityPrompts/RTP-LX 30,300
+AyaRedTeaming aya-redteaming-dataset CohereLabs/aya_redteaming 2662
+Thai Toxicity tweets thai-toxicity-tweets tmu-nlp/thai_toxicity_tweet 3,300
+Ukr Toxicity ukr-toxicity-dataset ukr-detect/ukr-toxicity-dataset 5,000
+HarmBench (HB) (mazeika2024-harmbench) walledai/HarmBench 400
+Forbidden Questions (FQ) (forbidden-questions-dataset) TrustAIRLab/forbidden_question_set 390
+Simple Safety Tests (simple-safety-tests) walledai/SimpleSafetyTests 100
+SaladBench (SaladB) (salad-bench-dataset) walledai/SaladBench 26,500
+Toxicity Jigsaw (TJS) toxicity-jigsaw Arsive/toxicity_classification_jigsaw 26,000
+Toxic Text (toxic-text-dataset) nicholasKluge/toxic-text 41,800
+AdvBench (advbench-dataset) walledai/AdvBench 520
+CodeAttack (ren-2024-codeattack-jailbreak) https://github.com/AI45Lab/CodeAttack 3120
+Vision JailBreakV-28K (luo2024jailbreakv28k) JailbreakV-28K/JailBreakV-28k 8,000
+VLSafe (chen2024-vlsafe-dataset) YangyiYY/LVLM_NLF 1,110
+FigStep (figstep-dataset-image-jailbreak) https://github.com/wangyu-ovo/MML 500
+MML SafeBench (mml-safebench-image-jailbreak) https://github.com/wangyu-ovo/MML 2,510
+Hades (hades-dataset-image-jailbreak) Monosail/HADES 750
+SafeBench (another-safebench-dataset-image-jailbreaks) Zonghao2025/safebench 2,300
+MM SafetyBench ($m$-safetybench-image-jailbreak) PKU-Alignment/$M$-SafetyBench 1680
+RedTeamVLM (li2024-RedTeamVLM-dataset) MMInstruction/RedTeamingVLM 200
+VLSBench (vlsbench-dataset-image-jailbreak) Foreshhh/vlsbench 2,240
+Audio VoiceBench (Alpacaeval) advbench-audio-jailbreaks hlt-lab/voicebench 636
+AIAH (achilles-dataset-audio-jailbreaks) https://github.com/YangHao97/RedteamAudioLMMs 350
 Table 1: Details of datasets used for training and evaluation.
@@
 The remaining text datasets are English-only, and were translated to 72 other languages (52 natural and 20 cipher): HarmBench (HB), Forbidden Questions (FQ), Simple Safety Tests, SaladBench (SaladB), Toxicity Jigsaw (TJS), Toxic Text, and AdvBench.
-Model name Citation HuggingFace ID Rough Parameter Count
-General Llama3.3-70B-Instruct (Grattafiori et al., 2024) meta-llama/Llama-3.3-70B-Instruct 70B
-Molmo-7B (Deitke et al., 2024) allenai/Molmo-7B-D-0924 7B
-LLaMA-Omni 8B (Fang et al., 2025) ICTNLP/Llama-3.1-8B-Omni 8B
-Kokoro (Hexgrad, 2025) hexgrad/Kokoro-82M 82M
-Text LlamaGuard 1 Inan et al. (2023) meta-llama/LlamaGuard-7b 7B
-LlamaGuard 2 Inan et al. (2023) meta-llama/Meta-Llama-Guard-2-8B 8B
-LlamaGuard 3 Inan et al. (2023) meta-llama/Llama-Guard-3-8B 8B
-AegisGuard Permissive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Permissive-1.0 7B
-AegisGuard Defensive Ghosh et al. (2024a) nvidia/Aegis-AI-Content-Safety-LlamaGuard-Defensive-1.0 7B
-WildGuard Han et al. (2024) allenai/wildguard 7B
-HarmBench (mistral) Mazeika et al. (2024) cais/HarmBench-Mistral-7b-val-cls 7B
-HarmBench (llama) Mazeika et al. (2024) cais/HarmBench-Llama-2-13b-cls 13B
-DuoGuard Deng et al. (2025) DuoGuard/DuoGuard-1B-Llama-3.2-transfer 1B
-PolyGuard Kumar et al. (2025) ToxicityPrompts/PolyGuard-Qwen 7B
-Vision Llama Guard 3 Vision (Chi et al., 2024) meta-llama/Llama-Guard-3-11B-Vision 11B
-VLMGuard (Du et al., 2024) --- 2.2M
-LLavaGuard (Helff et al., 2025) AIML-TUDA/LlavaGuard-7B-hf 7B
-Table 2: Model and baseline details.
 3.1 Selecting universal layers via the U-Score
@@
 To select language-agnostic layers, we use a dataset of translated sentence pairs spanning various languages.
-Specifically, we use sentences in 53 natural languages from the Flores200 dataset and additionally translate the sentences into 20 cipher languages (using encodings such as Caesar shifts, base64, hexadecimal); see a full list in Appendix A.
+Specifically, we use sentences in 53 natural languages from the Flores200 dataset and additionally translate the sentences into 20 cipher languages (using encodings such as Caesar shifts, base64, hexadecimal); see a full list in Appendix˜A.
 We extract embeddings from each layer of Llama3.3-70B-Instruct for the sentences in all 73 languages and use them to compute the U-Score (averaged over languages).
-Fig.
-2 shows the U-Score as a function of layer depth.
+Fig.˜2 shows the U-Score as a function of layer depth.
 For natural languages (blue curve), the U-Score peaks in the middle layers of the model, with the highest values in layer 57 (out of 81 layers).
-For cipher languages (red curve), the U-Score is much lower than for natural languages, suggesting the model fails to represent semantic similarity in these languages (see analysis in Section 5).
+For cipher languages (red curve), the U-Score is much lower than for natural languages, suggesting the model fails to represent semantic similarity in these languages (see analysis in Section˜5).
 Selecting modality-agnostic layers
 To select layers aligned between images and captions, we use the MM-Vet v2 dataset, a popular dataset for MLLM evaluation containing 517 examples, each consisting of a text question paired with one or more images.
-We generate captions for each image using a captioning model (Molmo-7B) and then extract embeddings for each image and its corresponding caption using an MLLM (also Molmo-7B) and use them to compute the U-Score, which peaks in layer 22 (out of 28 layers; see Fig.
-2 green curve).
+We generate captions for each image using a captioning model (Molmo-7B) and then extract embeddings for each image and its corresponding caption using an MLLM (also Molmo-7B) and use them to compute the U-Score, which peaks in layer 22 (out of 28 layers; see Fig.˜2 green curve).
 To select layers aligned between text and audio, we use the audio version of the Alpacaeval dataset from VoiceBench, a dataset of 636 audio-transcript pairs.
-We extract embeddings from each layer of an MLLM (LLaMA-Omni 8B) and use them to compute the U-Score, which peaks at layer 21 (out of 32 layers; see Fig.
-2 purple curve).
+We extract embeddings from each layer of an MLLM (LLaMA-Omni 8B) and use them to compute the U-Score, which peaks at layer 21 (out of 32 layers; see Fig.˜2 purple curve).
 Overall, we see that LLMs and MLLMs generate representations that are shared across languages and modalities.
@@
 OmniGuard classifier.
-Following Section 3.1, we build a classifier that takes as input embeddings from layer 57 of Llama3.3-70B-Instruct.
-As training data, we randomly select 2,800 examples from the Aegis AI Content Safety dataset, balancing the benign and harmful classes.
-Notably, this dataset is about 18$\times$ smaller than the training data used by our baseline methods.
-We translate these English examples to 52 other natural languages (via the Google Translate API) and 20 cipher languages (using fixed rules), totaling 73 languages.
-We train OmniGuard using only half the languages (see list in Appendix A).
+[…]
 Baselines.
-We compare to many popular guard models (see Table 2) middle row.
+We compare to many popular guard models (see Table˜2) middle row.
 Notably, DuoGuard and PolyGuard were trained to detect harmful prompts across multiple languages.
-For a more direct comparison, we also compare to finetuned versions of DuoGuard and PolyGuard using the same 37 languages we use to train OmniGuard; Following the original PolyGuard paper, we finetuned these models using LoRA (Hu et al., 2021) for all linear layers with rank 8 and alpha 16 for one epoch with a learning rate of $42e-$ 4.
+For a more direct comparison, we also compare to finetuned versions of DuoGuard and PolyGuard using the same 37 languages we use to train OmniGuard; Following the original PolyGuard paper, we finetuned these models using LoRA (lora-paper) for all linear layers with rank 8 and alpha 16 for one epoch with a learning rate of 2e42e-4.
 Datasets.
-We evaluate on several common text attack benchmarks (see Table 1).
+We evaluate on several common text attack benchmarks (see Table˜1).
 We additionally evaluate on three benchmarks from CodeAttacks that transform a harmful query as a list, a stack, or as a string in a Python program, obfuscating the harmfulness.
-For evaluation in this setup, we transform the harmful prompts from AdvBench and benign prompts from Toxicity Jigsaw datasets in the three code formats and subsampl the Toxicity Jigsaw dataset to be of the same size as Advbench.
-Note that, for this experiment, we only trained OmniGuard on the English subset of the training dataset.
+For evaluation in this setup, we transform the harmful prompts from AdvBench and benign prompts from Toxicity Jigsaw datasets in the three code formats and subsample the Toxicity Jigsaw dataset to be of the same size as Advbench.
+Note that for this experiment, we only trained OmniGuard on the English subset of the training dataset.
+Model name Citation HuggingFace ID Rough Parameter Count
+General Llama3.3-70B-Instruct (grattafiori2024llama) meta-llama/Llama-3.3-70B-Instruct 70B
+Molmo-7B (deitke2024-molmo-model) allenai/Molmo-7B-D-0924 7B
+LLaMA-Omni 8B (fang2025-llama-omni-model) ICTNLP/Llama-3.1-8B-Omni 8B
+Kokoro (hexgrad_kokoro_82m_2025) hexgrad/Kokoro-82M 82M
+Text LlamaGuard 1 inan2023-llama-guardrails meta-llama/LlamaGuard-7b 7B
+LlamaGuard 2 inan2023-llama-guardrails meta-llama/Meta-Llama-Guard-2-8B 8B
+LlamaGuard 3 inan2023-llama-guardrails meta-llama/Llama-Guard-3-8B 8B
+AegisGuard Permissive ghosh2024-aegis-safety-guard nvidia/Aegis-AI-Content-Safety-LlamaGuard-Permissive-1.0 7B
+AegisGuard Defensive ghosh2024-aegis-safety-guard nvidia/Aegis-AI-Content-Safety-LlamaGuard-Defensive-1.0 7B
+WildGuard Han2024-wildguardmix-dataset allenai/wildguard 7B
+HarmBench (mistral) mazeika2024-harmbench cais/HarmBench-Mistral-7b-val-cls 7B
+HarmBench (llama) mazeika2024-harmbench cais/HarmBench-Llama-2-13b-cls 13B
+DuoGuard duo-guard-multilingual DuoGuard/DuoGuard-1B-Llama-3.2-transfer 1B
+PolyGuard polyguard-model ToxicityPrompts/PolyGuard-Qwen 7B
+Vision Llama Guard 3 Vision (chi2024-llamaguard3-vision) meta-llama/Llama-Guard-3-11B-Vision 11B
+VLMGuard (du2024-vlmguard) --- 2.2M
+LLavaGuard (helff2025-llavaguard) AIML-TUDA/LlavaGuard-7B-hf 7B
+Table 2: Model and baseline details.
 3.2.2 Setup for vision attacks
 OmniGuard classifier.
-Following Section 3.1, we build a classifier that takes as input embeddings from layer 22 of Molmo-7B.
+Following Section˜3.1, we build a classifier that takes as input embeddings from layer 22 of Molmo-7B.
 As training data, we use 2000 image-query pairs randomly sampled from the JailBreakV-28K dataset and 1024 image-query pairs sampled from the VLSafe dataset as the harmful datapoints and 517 image-query from the MM-Vet v2 dataset as the benign datapoints.
 Baselines.
-We compare to guard models that take an image or image-text pair and output a binary harmfulness classification (see Table 2 bottom row). We train VLMGuard on the same training data as OmniGuard.
+We compare to guard models that take an image or image-text pair and output a binary harmfulness classification (see Table˜2 bottom row). We train VLMGuard on the same training data as OmniGuard.
 Datasets.
-We evaluate detecting image/text attacks using several datasets (see Table 1).
+We evaluate detecting image/text attacks using several datasets (see Table˜1).
 FigStep and MML Safebench are typographic attacks that embed a harmful prompt in an image.
@@
 MM-safetybench, RTVLM, and VLSBench consist of an image and a query where the text query is seemingly benign, but when combined with the respective image, it is harmful (e.g.
-see Figure 1).
+see Figure˜1).
 3.2.3 Setup for audio attacks
 OmniGuard classifier.
-Following Section 3.1, we build a classifier that takes as input embeddings from layer 21 of Llama-Omni-8B.
+Following Section˜3.1, we build a classifier that takes as input embeddings from layer 21 of Llama-Omni-8B.
 We train the classifier on the English portion of the training data we use for the text setting, by using a text-to-speech model to convert the text into audio.
@@
 We are unaware of any existing models for detecting harmful audio input.
-The most relevant approach, SpeechGuard (Peri et al., 2024) adds noise as a defense against potentially harmful audio inputs but does not directly classify harmfulness.
+The most relevant approach, SpeechGuard (peri-etal-2024-speechguard) adds noise as a defense against potentially harmful audio inputs but does not directly classify harmfulness.
 To contextualize our results for audio benchmarks, we compare performance to guard models that directly classify the raw text present in the audio (OmniGuard and LlamaGuard3).
 Datasets.
-We use the two audio benchmarks (see Table 1 bottom row).
+We use the two audio benchmarks (see Table˜1 bottom row).
 We also evaluate on several text jailbreak benchmarks using Kokoro to convert them from text to speech: HB, FQ, Simple Safety Tests, SaladB, and TJS.
@@
 In all settings, OmniGuard achieves the highest performance.
-Table B1 further stratifies these results by high-resource, low-resource, and cipher languages.
+Table˜B1 further stratifies these results by high-resource, low-resource, and cipher languages.
 Hades VLSBench MM-SafetyBench SafeBench RTVLM FigStep Avg.
@@
 Defending against multilingual text attacks
-Table 3 compares the accuracy of detecting harmful prompts for text benchmarks.
-Table 3(A) shows results for multilingual benchmarks, where OmniGuard achieves the highest accuracy (86.36%) compared to the baselines, and achieves new state-of-the-art performance for 3 benchmarks: MultiJail, RTP-LX, and AyaRedTeaming.
+Table˜3 compares the accuracy of detecting harmful prompts for text benchmarks.
+Table˜3(A) shows results for multilingual benchmarks, where OmniGuard achieves the highest accuracy (86.36%) compared to the baselines, and achieves new state-of-the-art performance for 3 benchmarks: MultiJail, RTP-LX, and AyaRedTeaming.
 The strongest baseline is Polyguard, which yields an average accuracy of 83.19%, despite being trained on a much larger dataset (1.91M examples for Polyguard versus 103K examples for OmniGuard).
-In benchmarks that were translated from English to various other languages, including cipher languages, we again see that OmniGuard achieves the highest accuracy (Table 3(B)).
-Finally, Table 3(C) shows that OmniGuard outperforms finetuned versions of DuoGuard and Polyguard on unseen languages, demonstrating that OmniGuard can outperform methods that were trained specifically for multilingual harmfulness classification.
+In benchmarks that were translated from English to various other languages, including cipher languages, we again see that OmniGuard achieves the highest accuracy (Table˜3(B)).
+Finally, Table˜3(C) shows that OmniGuard outperforms finetuned versions of DuoGuard and Polyguard on unseen languages, demonstrating that OmniGuard can outperform methods that were trained specifically for multilingual harmfulness classification.
 Defending against image-based attacks
-Table 4 shows the accuracy of detecting harmful image and text prompts for (A) pairs consisting of images and text queries, where either the image or both the image and query can be harmful and (B) typographic images with various encryptions.
+Table˜4 shows the accuracy of detecting harmful image and text prompts for (A) pairs consisting of images and text queries, where either the image or both the image and query can be harmful and (B) typographic images with various encryptions.
 OmniGuard achieves the highest performance for both sets of benchmarks (95.44% and 79.76%) while being trained using only about 3500 image-query pairs (compared to about 5500 datapoints used by LlavaGuard).
@@
 Defending against audio-based attacks
-Table 5 shows the accuracy of detecting harmful audio prompts.
+Table˜5 shows the accuracy of detecting harmful audio prompts.
 OmniGuard detects harmful audio input with high accuracy across all benchmarks.
@@
 We also evaluate the accuracy of OmniGuard and baselines in adapting to out-of-distribution code attacks given very few samples.
-In this setting, some prior work has speculated that guard models may be very data efficient, as they can make use of few-shot examples in-context (Inan et al., 2023).
-However, we find that baseline guard models generally struggle to rapidly adapt to this setting given few-shot examples (Figure 3).
-[^2] Note, we omit baseline guard models that achieve 90% accuracy or greater without any few-shot examples, as their training data likely explicitly includes code attacks.
+In this setting, some prior work has speculated that guard models may be very data efficient, as they can make use of few-shot examples in-context (inan2023-llama-guardrails).
+However, we find that baseline guard models generally struggle to rapidly adapt to this setting given few-shot examples (Figure˜3).
+[^2] Note that we omit baseline guard models that achieve 90% accuracy or greater without any few-shot examples, as their training data likely explicitly includes code attacks.
 In contrast, OmniGuard is able to rapidly achieve close to 100% accuracy for all three benchmarks by updating its lightweight parameters using less than five examples.
@@
 For the text-only model, we compare the U-Score-selected layer (57) to 3 other layers (layer 10, layer 75, and the last layer) when used for a set of toxicity prediction tasks.
-Table 6 shows that the representations from the layer with the highest U-Score result in significantly better harmfulness classification accuracy, improving between 5% and 14% compared to the other layers.
+Table˜6 shows that the representations from the layer with the highest U-Score result in significantly better harmfulness classification accuracy, improving between 5% and 14% compared to the other layers.
+We show ablation over more layers in Table˜B2.
 Thai Tox Ukr Tox TJS ToxText Avg.
@@
 U-Score selected layer (57) 68.7 73.1 76.8 75.07 73.4
-Table 6: OmniGuard’s accuracy of detecting harmful prompt when trained using representation from different model layers.
+Table 6: OmniGuard’s accuracy of detecting harmful prompts when trained using representations from different model layers.
 OmniGuard’s efficiency
-OmniGuard is highly efficient at inference time because it re-uses the internal representations of the main LLM that is already processing the user query for generation.
-Therefore, its compute time is only that of a lightweight multilayer perceptron, making it much faster than baseline guard models (note that this does limit OmniGuard to only work when the generation model is open-source, so embeddings can be extracted).
-Table 7 shows the inference time required by various guard models to predict the harmfulness of prompts in the AdvBench dataset in English, translated to Spanish, French, Telugu, and base64 encoding.
-OmniGuard is the fastest and is about 120$\times$ faster than the fastest baseline (DuoGuard).
-Inference time as measured on a machine with 1 L40 GPU, 4 CPUs, and 50 GB RAM.
+[…]
 Guard Method Inference Time (s) $\downarrow$
@@
 OmniGuard 0.04
-Table 7: Average inference time required for harmfulness prediction on the AdvBench dataset (averaged over 5 languages). OmniGuard is about 120$\times$ faster than the fastest baseline (DuoGuard).
+[…]
+Performance comparison across base LLMs
+We compare OmniGuard’s accuracy when using different base LLMs in Table˜B3.
+We trained the classifiers on the layers with the best U-scores for each model.
+We find that the average accuracy for the moderator model trained using smaller LLMs is lower than the moderator model trained using the larger Llama3.3-70B-Instruct model.
 Performance comparison across languages.
-We now analyze the harmfulness classification accuracy of OmniGuard by language, and compare it to the underlying LLM’s sentiment classification accuracy for the same language (Fig.
-4).
-We measure harmfulness classification accuracy using OmniGuard on all the datasets in Table 3 and sentiment classification accuracy using Llama3.3-70B-Instruct with zero-shot prompting on 72 translated versions of the SST-2 dataset (translated to all the languages we consider).
-We observe that the accuracies are generally correlated indicating that OmniGuard is able to defend well for languages for which the LLM is more coherent / susceptible to attack.
+We now analyze the harmfulness classification accuracy of OmniGuard by language, and compare it to the underlying LLM’s sentiment classification accuracy for the same language (Fig.˜4).
+We measure harmfulness classification accuracy using OmniGuard on all the datasets in Table˜3 and sentiment classification accuracy using Llama3.3-70B-Instruct with zero-shot prompting on 72 translated versions of the SST-2 dataset (translated to all the languages we consider).
+We observe that the accuracies are generally correlated, indicating that OmniGuard is able to defend well in languages for which the LLM is more coherent/susceptible to attack.
 Unsurprisingly, the accuracies for natural languages are higher than the accuracies for cipher languages.
@@
 Several techniques have recently emerged to attack or jailbreak LLMs.
-Early techniques relied on manual effort and were very time-intensive Shen et al.
-(2024b); Andriushchenko et al.
-(2024).
-Later techniques automated this process, e.g., Zou et al.
-(2023b); Jones et al.
-(2023); Zhu et al.
-(2023) proposed gradient-based approaches to identify inputs to jailbreak LLMs with white-box access.
-Another set of techniques start from a set of human written prompts and modify them using approaches like genetic algorithms (Liu et al., 2024a; Lapid et al., 2024; Li et al., 2024d), fuzzing (Yu et al., 2024a), or reinforcement learning (Chen et al., 2024b) to automatically produce prompts for jailbreaking.
-Another set of techniques, use a helper LLM to generate prompts that attack a target LLM (Chao et al., 2024; Ding et al., 2024; Mehrotra et al., 2024).
-Finally, Wei et al.
-(2024); Wang et al.
-(2023); Anil et al.
-(2024); Pernisi et al.
-(2024) use simple in-context demonstrations to jailbreak the models by overcoming its safety training and Russinovich et al.
-(2024); Li et al.
-(2024c) propose using multi-turn dialogues to jailbreak models.
+Early techniques relied on manual effort and were very time-intensive shen2024-do-now-evaluating; andriushchenko2024-adaptive-attacks-best-asr-paper.
+Later techniques automated this process, e.g., zou2023-gcg-attacks; jones-another-gcg-attack; zhu2023-autodan-interpretable-gradient proposed gradient-based approaches to identify inputs to jailbreak LLMs with white-box access.
+Another set of techniques start from a set of human written prompts and modify them using approaches like genetic algorithms (liu2024-autodan; lapid2024-opensesame-another-genetic; li2024-semanticmirror-another-genetic), fuzzing (yu2024-gpt-fuzzer-jailbreak), or reinforcement learning (chen2024-rl-jailbreak) to automatically produce prompts for jailbreaking.
+Another set of techniques, use a helper LLM to generate prompts that attack a target LLM (chao2024-PAIR-method; ding-2024-nested-attacks; mehrotra2024-tree-attacks-jailbreak).
+Finally, wei2024-icl-jailbreak1; wang2023-icl-jailbreak2; anthropic-many-shot-jailbreak; pernisi2024-many-shot-jailbreak-italian use simple in-context demonstrations to jailbreak the models by overcoming its safety training and russinovich2024-multi-turn-crescendo-attack; scale-ai-multi-turn-dataset propose using multi-turn dialogues to jailbreak models.
 Multilingual Jailbreak Attacks
@@
 To tackle this, a novel set of techniques have emerged that attack models using inputs in various languages or obfuscations that are able to bypass the safety guardrails.
-Deng et al.
-(2024); Yong et al.
-(2024); Wang et al.
-(2024a); Yang et al.
-(2024); Yoo et al.
-(2024); Upadhayay and Behzadan (2024); Song et al.
-(2024) demonstrated that attacking models using mid and low resources languages led to higher attack success rates, compared to the case of attacking the model in high-resource languages like English.
-Going beyond natural languages, a newer set of works propose using cipher characters or languages to evade the safety filters, e.g., Jin et al.
-(2024) propose interspersing cipher characters in between text, Jiang et al.
-(2024) propose replacing the unsafe words with their ASCII art versions, and Yuan et al.
-(2024a) propose prompting models in cipher languages like Morse, Atbash, Caesar.
+deng2024-multilingual-attack-firstpaper; yong2024-low-resource-languages-jailbreak-gpt4; wang2024-multilingualsafety; dan-roth-paper-guard-llms-on-multilingual-data; yoo2024-switch-between-languages; upadhayay2024-sandwich-attack-multilanguagemixture; song2024-multilingual-attacks demonstrated that attacking models using mid and low resources languages led to higher attack success rates, compared to the case of attacking the model in high-resource languages like English.
+Going beyond natural languages, a newer set of works propose using cipher characters or languages to evade the safety filters, e.g., jin2024-JAMBench propose interspersing cipher characters in between text, jiang2024-ascii-jailbreak propose replacing the unsafe words with their ASCII art versions, and yuan2024-gpt-cipher-attack propose prompting models in cipher languages like Morse, Atbash, Caesar.
 Multimodal Jailbreak Attacks
@@
 Several recent works have shown that MLLMs remain vulnerable to being jailbroken when prompted with images or audios that have a harmful query (the same harmful query in text would be easily detected as harmful).
-Liu et al.
-(2024b) show that using a prompt with a correlated image, e.g., using an image of a bomb when asking the model to answer the question: How to make a bomb?
+mm-safetybench-image-jailbreak show that using a prompt with a correlated image, e.g., using an image of a bomb when asking the model to answer the question: How to make a bomb?
 is more likely to jailbreak a model than when using an uncorrelated image.
-Hu et al.
-(2025) argue that providing a harmful image with a benign query (see Figure 1) further increases the potential of jailbreaking the model.
-Gong et al.
-(2025) and Wang et al.
-(2024b) demonstrate jailbreaking models by simply typographically embedding harmful queries in an image.
+vlsbench-dataset-image-jailbreak argue that providing a harmful image with a benign query (see Figure˜1) further increases the potential of jailbreaking the model.
+figstep-dataset-image-jailbreak and mml-safebench-image-jailbreak demonstrate jailbreaking models by simply typographically embedding harmful queries in an image.
 Safety moderation in LLMs
 Safety moderation in LLMs broadly fits into two categories: intrinsic and extrinsic.
-Intrinsic mechanisms include finetuning or RLHF training on an LLM (Bianchi et al., 2024; Chen et al., 2024a; Yuan et al., 2024b; Ouyang et al., 2022b; Bai et al., 2022; Dai et al., 2023).
+Intrinsic mechanisms include finetuning or RLHF training on an LLM (bianchi2024-finetuning-for-safety; chen2024-finetuning-for-safety; yuan2024-refuse-feel-unsafe; rlhf-paper1; rlhf-paper2; dai2023-safe-rlhf).
 Extrinsic safety mechanisms utilize external models to detect harmful inputs and responses; these models can either be simple filters or use guard models.
-Jain et al.
-(2023); Alon and Kamfonas (2023); Hu et al.
-(2024) propose using perplexity filtering for detecting harmful prompts.
-Guard models defend LLMs by training separate LLMs to detect harmful text (see Table 2).
+jain2023-perplexity-defense1; alon2023-perplexity-defense2; hu2024-perplexity-defense3 propose using perplexity filtering for detecting harmful prompts.
+Guard models defend LLMs by training separate LLMs to detect harmful text (see Table˜2).
+Separately, a line of work has introduced interpretability methods to transparently expose safety concerns in LLMs (bereska2024mechanistic; singh2023augmenting; arditi2024refusal; benara2024crafting).
+A few other works also use internal model representations to defend against harmful inputs.
+However, they defer from our work: OmniGuard is a standalone content safety classification model while the other approaches like Jailbreak Antidote (shen2025jailbreak-antidote) and AdaSteer (zhao2025-ada-steer) directly change the internal representations to defend against harmful inputs.
+Therefore, OmniGuard is similar to other safety classification models like LlamaGuard and WildGuard.
+Additionally, OmniGuard can detect harmful prompts in multiple languages (both natural and cipher) and multiple modalities using the same method while the other approaches only defend against attacks in English text.
+OmniGuard is also extremely efficient in producing safety predictions while the other approaches take as much time as the inference of the underlying LLM, which can typically take several seconds to minutes.
 Safety against multilingual attacks
-DuoGuard (Deng et al., 2025) and PolyGuard (Kumar et al., 2025) are the two previous guard models that were specifically trained to defend against multilingual attacks.
+DuoGuard (duo-guard-multilingual) and PolyGuard (polyguard-model) are the two previous guard models that were specifically trained to defend against multilingual attacks.
 DuoGuard uses a two-player RL-driven mechanism to generate harmful data in multiple languages and uses that to finetune a Llama3.2-1B model.
@@
 Safety moderation in MLLMs
-Relatively few works have tackled detecting harmful prompts in multimodal settings (see Table 1 and Table 2).
-Chi et al.
-(2024) propose LlamaGuard3-11B-Vision (a finetuned version of Llama-3-11B-Vision) for detecting unsafe inputs in images and texts.
-Du et al.
-(2024) and Helff et al.
-(2025) propose other approaches for the same task.
+Relatively few works have tackled detecting harmful prompts in multimodal settings (see Table˜1 and Table˜2).
+chi2024-llamaguard3-vision propose LlamaGuard3-11B-Vision (a finetuned version of Llama-3-11B-Vision) for detecting unsafe inputs in images and texts.
+du2024-vlmguard and helff2025-llavaguard propose other approaches for the same task.
 OmniGuard achieves higher accuracy in detecting harmful images and prompts compared to these approaches, and to the best of our knowledge is the first guard model for harmful audio inputs.
@@
 Table B1: Accuracy of detecting harmful prompts stratified by high-resource natural, low-resource natural, and cipher languages.
+Appendix C Experimental Details of Filtering of Wikitext and its Translation
+Thai Tox Ukr Tox TJS ToxText Avg.
+Layer 10 62.1 65.5 66.95 61.89 64.42
+Layer 55 67.4 73.0 76.91 74.96 73.06
+Layer 56 66.8 71.5 74.54 71.82 71.17
+Selected layer 57 68.7 73.1 76.8 75.07 73.40
+Layer 58 66.6 73.2 74.92 72.63 71.84
+Layer 59 67.3 73.3 76.44 74.22 72.82
+Layer 60 67.5 72.6 76.43 74.46 72.75
+Layer 61 67.8 70.9 74.77 72.76 71.56
+Layer 62 66.1 72.3 74.78 72.83 71.50
+Layer 75 65.2 66.4 70.72 65.79 68.26
+Last Layer 63.1 51.2 61.33 56.76 59.05
+Table B2: OmniGuard’s accuracy of detecting harmful prompts when trained using representations from different model layers.
+Appendix D OmniGuard’s Performance With Different Base LLMs
+Models HarmBench FQ SimpleST SaladB TJS ToxText AdvBench Avg.
+Llama3.1-8B-Instruct 81.16 88.72 88.78 87.35 71.32 69.20 88.62 82.16
+Gemma-3-4B-Instruct 78.03 88.23 93.86 82.09 53.50 51.45 89.07 76.60
+Qwen-3-4B-Instruct 82.06 86.25 80.99 82.15 58.83 57.70 84.80 76.11
+Olmo-2-7B-Instruct 80.57 88.86 92.14 88.68 70.65 64.79 88.24 81.99
+Mistral-8B-Instruct 84.27 87.23 89.82 88.08 72.49 68.71 89.07 82.81
+Llama3.3-70B-Instruct 89.13 89.57 89.62 87.30 76.68 75.07 86.59 84.85
+Table B3: OmniGuard’s accuracy of detecting harmful prompts when paired with different underlying LLMs across multiple benchmarks.